### v2

# Integrable sub-Riemannian geodesic flows on the special orthogonal group

Abstract.
We analyse the geometry of the rolling distribution on the special orthogonal group and show that almost all right-invariant sub-Riemannian metrics defined on this distribution have completely integrable geodesic flows. Our argument is an adaptation of the method used to establish integrability of the Riemannian metric arising from the $n$-dimensional rigid body: namely, by exhibiting a Lax pair and bi-Hamiltonian structure for the reduced equations of motion.
1. Introduction
The geodesic flow on a Riemannian manifold $M$ is generated by a Hamiltonian system on the cotangent bundle. A Riemannian metric is called integrable if the corresponding Hamiltonian system is integrable in the sense of Liouville; that is, if the the flow admits $n=\dim M$ independent integrals of motion which mutually commute with respect to the Poisson bracket on $T^{*}M$. In principle, one can then find the geodesics explicitly through quadrature. The literature surrounding integrable Riemannian metrics is vast (see for instance [6, 33, 29, 11, 10]) whereas less has been said about sub-Riemannian metrics, with much of the work focussed on Carnot groups [22, 25, 12]. A sub-Riemannian metric also defines a Hamiltonian system on the cotangent bundle, and so one can equally ask if this flow is integrable.
A famous example of an integrable Riemannian metric on $\mathbf{SO}(n)$ is that which arises from the kinetic energy of an $n$-dimensional rigid body. We will consider a related class of sub-Riemannian metrics defined on the right-invariant ‘rolling’ distribution $\mathcal{D}$ of tangent vectors to $\mathbf{SO}(n)$ of the form $\dot{g}=\Omega g$ where
(1) $\Omega=\left(\begin{array}[]{c|c}0&\begin{matrix}\Omega_{12}&\dots&\Omega_{1n}\end{matrix}\\ \hline\cr\begin{matrix}-\Omega_{12}\\ \vdots\\ -\Omega_{1n}\end{matrix}&\mbox{\huge$0$}\end{array}\right)$
are skew-symmetric matrices whose only non-zero entries are in the first row and column.
Theorem 1.
The sub-Riemannian metric with length
(2) $\sqrt{A_{2}\Omega_{12}^{2}+\dots+A_{n}\Omega_{1n}^{2}}$
defined on the distribution $\mathcal{D}$ defines an integrable Hamiltonian system on $T^{*}\mathbf{SO}(n)$ for any choice of distinct positive scalars $A_{2},\dots,A_{n}$.
We prove this theorem by adapting the same methods which led Manakov to establish the complete integrability of the rigid body [23]. In particular, we consider the symplectic reduced space $\mathfrak{so}(n)^{*}$ and show that the corresponding reduced equations of motion admit a Lax pair formulation and a bi-Hamiltonian structure. This yields a collection of first integrals which can then be shown to define an integrable system thanks to the work of Mishchenko & Fomenko in [24].
Any sub-Riemannian problem may be reinterpreted as a kinematic non-holonomic optimal control problem. From this point of view, our problem is a special case of a more general class of problems defined on symmetric spaces considered in [7, 4]. There is also a wider literature concerning the related problem of rolling manifolds along each other, see for instance [19, 17]. In particular, we note that Theorem 1 was proven in the symmetric case where $A_{2}=\dots=A_{n}$ for $n=3$ in [3] and more generally in [20] where the sub-Riemannian problem corresponds to a ball rolling on a plane.
2. A sub-Riemannian metric on the special orthogonal group
2.1. The Hamiltonian setting
A sub-Riemannian metric on a manifold $M$ is completely determined by a fibrewise self-adjoint map $\beta\colon T^{*}M\rightarrow TM$ of constant rank known as the cometric. The distribution $\mathcal{D}$ is the image of $\beta$ and the inner product between $\beta(p_{1})$ and $\beta(p_{2})$ is defined as $\langle p_{1},\beta(p_{2})\rangle$. The function
$H(p)=\frac{1}{2}\langle p,\beta(p)\rangle$
on $T^{*}M$ is the Hamiltonian for the sub-Riemannian metric. It is known that the integral curves of the Hamiltonian flow when projected down to $M$ are, for sufficiently small arcs, the unique length-minimising admissible curves between their endpoints [26].
We shall be interested in the case when the manifold is a Lie group $G$. Left and right multiplication of the group on itself lifts to an action on the cotangent bundle, and we will say that the sub-Riemannian metric on $G$ is left- or right-invariant if the Hamiltonian is invariant with respect to the left or right action.
The momentum maps
${T^{*}G\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathbf{J}_{L}}$$\scriptstyle{\mathbf{J}_{R}}$${{\mathfrak{g}^{*}_{+}}}$${{\mathfrak{g}^{*}_{-}}}$
for the left and right actions are given by right and left translation of covectors back to the identity, respectively. For left-invariant Hamiltonians the momentum $\mathbf{J}_{L}$ is constant along the Hamiltonian flow, and $\mathbf{J}_{R}$ serves as an orbit-quotient; and vice versa with the words left and right interchanged. The orbit space $\mathfrak{g}_{\pm}^{*}$ is Poisson with the bracket
(3) $\{f,g\}(M)=\pm\langle M,[df,dg]\rangle$
[…]
2.2. The rigid body
The kinetic energy for an $n$-dimensional rigid body defines a left-invariant Riemannian metric on $\mathbf{SO}(n)$. This descends through reduction to $\mathfrak{so}(n)^{*}_{-}$ with Hamiltonian
$H(M)=\frac{1}{2}\langle M,\Omega\rangle,\quad\text{for}\leavevmode\nobreak\ \Omega=\mathbb{J}^{-1}(M)$
where $\mathbb{J}\colon\mathfrak{so}(n)\rightarrow\mathfrak{so}(n)^{*}$ is the inertia operator [30]. If we identify $\mathfrak{so}(n)$ with its dual via the trace pairing $\langle M,\Omega\rangle=\operatorname{Tr}(M^{T}\Omega)$ then
$\mathbb{J}(\Omega)=J\Omega+\Omega J$
where $J$ is the mass matrix, which we may suppose is diagonal with strictly positive entries $J_{1},\dots,J_{n}$. The angular momentum is related to the angular velocity via
$M_{kl}=I_{kl}\Omega_{kl},\quad\text{for}\leavevmode\nobreak\ I_{kl}=J_{k}+J_{l}.$
The $I_{kl}$ are the generalised moments of inertia for the $n$-dimensional body, and the Hamiltonian is in coordinates given by
(4) $H(M)=\frac{1}{2}\sum_{k<l}\frac{M_{kl}^{2}}{I_{kl}}.$
The equations of motion on $\mathfrak{so}(n)^{*}_{-}$ are known as the Euler equations, and given by
$\frac{d}{dt}M=[M,\Omega].$
The trivial constants of motion are the Hamiltonian and the Casimirs $\operatorname{Tr}(M^{k})$. Additional constants of motion can be seen by rewriting the Euler equations in terms of a Lax pair
$\frac{d}{dt}(M+\lambda J^{2})=[M+\lambda J^{2},\Omega+\lambda J]$
where $\lambda$ is an arbitrary parameter. This equation implies that solution curves remain inside the same conjugacy class, and so $\operatorname{Tr}(M+\lambda J^{2})^{k}$ is constant for any power $k$. Since $\lambda$ was arbitrary it follows that each of the coefficients of $\lambda$ in the expansion
(5) $h^{\lambda}_{k}=\operatorname{Tr}(M+\lambda J^{2})^{k}=\sum_{l=0}^{k}\lambda^{k-l}h_{k,l}(M),$
is itself a constant of motion. These define the Manakov integrals $h_{k,l}$.
2.3. The sub-Riemannian metric
Let $\mathfrak{p}$ denote the subspace of matrices of the form (1) and equip this with the metric given in (2). We note that this is not a Lie subalgebra of $\mathfrak{so}(n)$, but that it generates the entire algebra in one step, that is
$\mathfrak{p}+[\mathfrak{p},\mathfrak{p}]=\mathfrak{g}.$
Left/right multiplication pushes $\mathfrak{p}$ forward to define a non-integrable left/right-invariant distribution on $\mathbf{SO}(n)$ equipped with a sub-Riemannian metric. The resulting length functional on the left-invariant distribution is unphysical in the sense that it is not the Lagrangian for a constrained physical body in $\mathbf{R}^{n}$. However, for the right-invariant distribution we can interpret the length functional as describing the length traced by the path of an $(n-1)$-dimensional ball rolling without slipping on $\mathbf{R}^{n-1}$ with metric
$ds^{2}=A_{2}dx_{2}^{2}+\dots+A_{n}dx_{n}^{2}.$
For this reason we shall work exclusively with the right-invariant distribution which we denote by $\mathcal{D}$ and refer to as the rolling distribution. We will later solve for the rolling motion in the case $n=3$.
The right-invariant sub-Riemannian Hamiltonian on $T^{*}\mathbf{SO}(n)$ descends to give a Lie-Poisson system on $\mathfrak{so}(n)^{*}_{+}$ whose Hamiltonian is given in coordinates by
(6) $H_{\text{sR}}(M)=\frac{1}{2}\sum_{k=2}^{n}\frac{M_{1k}^{2}}{A_{k}},$
where, as we did earlier, we have identified the Lie algebra with its dual by using the trace form.
Remark 2.1.
Notice that the sub-Riemannian Hamiltonian in (6) is the limit of the rigid-body Hamiltonian in (4) when $I_{kl}\rightarrow\infty$ for $k\neq 1$ and $A_{k}=I_{1k}$. Indeed, any sub-Riemannian metric on a distribution $\mathcal{D}$ can be interpreted as a limit of Riemannian penalty metrics in which velocity vectors not belonging to $\mathcal{D}$ become increasingly expensive. We note however that this limit is not physical since $2J_{1}=I_{12}+I_{13}-I_{23}$ necessarily limits to $\infty$- ∞ which contradicts the positive-definiteness of the mass matrix for realistic bodies. On the other hand, the metric on the left-invariant distribution does correspond to the optimal control of an underactuated $n$-dimensional rigid body, with control torques (actuators) about $(n-1)$ of the $n$ principal axes. For the underactuated rigid body problem see e.g. [7] and [4] and references therein.
2.4. Rolling a ball across a table
Place a ball on a table. The configuration of the ball is determined by its point of contact $(y,z)\in\mathbf{R}^{2}$ and a matrix $g\in\mathbf{SO}(3)$ whose columns are the basis vectors of an orthonormal frame fixed inside the ball. How can we roll the ball from one orientation $g_{0}$ to another $g_{1}$ in such a way that the curve of contact which the ball traces on the table has the least length with respect to the metric
$ds^{2}=A_{2}dy^{2}+A_{3}dz^{2}?$
This question is equivalent to our sub-Riemannian problem on $\mathbf{SO}(n)$ for $n=3$, as we will now show.
The rolling constraint implies that admissible tangent vectors to $\mathbf{SO}(3)$ must be of the form
(7) $\frac{dg}{dt}=\begin{pmatrix}0&-\Omega_{12}&\Omega_{13}\\ \Omega_{12}&0&0\\ -\Omega_{13}&0&0\end{pmatrix}g=\widehat{\Omega}g.$
This notation is chosen to coincide with the ‘hat-map’ whereby $\widehat{\Omega}v=\Omega\times v$ for $\Omega=(0,\Omega_{13},\Omega_{12})^{T}$. For such a motion, the path traced on the plane satisfies
(8) $\frac{d}{dt}\begin{pmatrix}y\\ z\end{pmatrix}=\begin{pmatrix}\Omega_{12}\\ -\Omega_{13}\end{pmatrix},$
where we are supposing for simplicity that the ball has radius $\pi^{-1}$. For an admissible curve $g(t)$ the length of the path traced by the ball is therefore
[…]
Eq. (7) defines a right-invariant distribution on $\mathbf{SO}(3)$ and Eq. (9) equips this with a sub-Riemannian metric.
[…]
Figure 1. On the right are intersections of the elliptic-cylinder $H_{\text{sR}}=1/2$ with coadjoint orbits in $\mathfrak{so}(3)^{*}$. The corresponding paths which the ball traces on the table are shown on the left.
Identify tangent vectors $\dot{g}$ to $\mathbf{SO}(3)$ with vectors $\Omega$ in $\mathbf{R}^{3}$ via right-trivialisation $\widehat{\Omega}=\dot{g}g^{-1}$ and the hat-map. Likewise, we may also identify covectors with $M$ in $\mathbf{R}^{3}$ via the standard pairing $\langle M,\Omega\rangle=M^{T}\Omega$. The Hamiltonian on $T^{*}\mathbf{SO}(3)$ for the sub-Riemannian metric is then
$H_{\text{sR}}(M)=\frac{1}{2}\left(\frac{M_{12}^{2}}{A_{2}}+\frac{M_{13}^{2}}{A_{3}}\right)$
for $M=(M_{23},M_{13},M_{12})^{T}$. After reduction this descends to a Lie-Poisson system in $\mathfrak{so}(3)^{*}_{+}\cong\mathbf{R}^{3}$ with equation of motion
(10) $\frac{d}{dt}M=\nabla H_{\text{sR}}\times M.$
The Hamiltonian $H_{\text{sR}}$ and Casimir function $|M|^{2}$ are both constants of motion. The intersections of their levels sets therefore give the solution curves: elliptic cylinders for $H_{\text{sR}}$ constant, and spheres of constant $|M|^{2}$, as shown on the right-hand side in Figure 1.
For the initial conditions $M_{12}(0)=\sqrt{C-A_{2}}$, $M_{12}(0)=\sqrt{A_{2}}$ and $M_{13}(0)=0$ we can solve (10) exactly in terms of Jacobi elliptic functions, finding
(11) $M_{12}(t)=\sqrt{A_{2}}\operatorname{cn}(u,k)\quad\text{and}\quad M_{13}(t)=\sqrt{A_{3}}\operatorname{sn}(u,k),$
where
$u=t\sqrt{\frac{C-A_{2}}{A_{2}A_{3}}}\quad\text{and}\quad k=\sqrt{\frac{A_{3}-A_{2}}{C-A_{2}}}.$
The tangent vector $\Omega$ is related to the covector $M$ via $M_{12}=A_{2}\Omega_{12}$ and $M_{13}=A_{2}\Omega_{13}$. We can use this relation together with (8) to find the path that the ball traces by integrating (11) (with the aid of [13]) to find that
(12) $\cos\left(z\sqrt{\textstyle\frac{A_{3}-A_{2}}{A_{3}}}\right)=\begin{cases}\sqrt{1-k^{2}}\cosh\left(\alpha-y\sqrt{\frac{A_{3}-A_{2}}{A_{2}}}\right),&0\leq k<1\\[20.0pt] \exp\left(-y\sqrt{\frac{A_{3}-A_{2}}{A_{2}}}\right),&k=1\\[20.0pt] \sqrt{k^{2}-1}\sinh\left(\alpha-y\sqrt{\frac{A_{3}-A_{2}}{A_{2}}}\right),&k>1\end{cases}$
where we are assuming $A_{3}>A_{2}$ and $\alpha$ is such that $y(0)=z(0)=0$. These curves are shown on the left-hand side in Figure 1.
Remark 2.2.
The rolling problem we have discussed is related to an interview question asked in the 1950s at the University of Oxford [16], first solved in [2] (see also [18]). Given two points $p_{0}$ and $p_{1}$ on the table, and two configurations $g_{0}$ and $g_{1}$ of the ball, what is the shortest length path which rolls $(p_{0},g_{1})$ to $(p_{1},g_{1})$? Our problem differs from this in two respects. Firstly, we do not care about connecting two points on the table, only the two configurations of the ball; and secondly, the metric on the table is not the standard Euclidean metric. It is curious to compare the curves in Eq. (12) with those for the Euclidean case $A_{2}=A_{3}$ where the ball is found to travel across the table in straight lines and circular arcs.
3. Integrability
3.1. Manakov integrals in the limit
Following on from Remark 2.1 we may interpret the sub-Riemannian metric as a limit of Riemannian metrics for the rigid body. Since the rigid body is known to be integrable for distinct principle moments of inertia, does it follow that its limit is also integrable? Our initial heuristic for finding first integrals was to inspect the Manakov integrals in such a limit. However, the integrals are found to diverge or become dependent. To overcome this, we now show how to define a collection of integrals which are functionally equivalent to the Manakov integrals, but which survive in the limit. We will later show that these are in fact first integrals of the sub-Riemannian flow.
To define this limit more precisely consider a rigid-body parametrised by $s$ whose mass matrix $J^{s}$ is diagonal with entries
$J_{1}^{s}=-s,\text{ and }J_{k}^{s}=A_{k}+s\text{ for }k\neq 1.$
In this way $I_{1k}=A_{k}$ for all $s$, and the $I_{kl}\rightarrow\infty$ as $s\rightarrow\infty$ for $k\neq 2$. The Hamiltonian therefore limits to the sub-Riemannian Hamiltonian $H_{\text{sR}}$. Strictly speaking this metric does not represent a physical body since $J^{s}_{1}$ is negative, however this does not affect the definition of the Manakov integrals. We may rewrite
$h_{k}^{\lambda}=\operatorname{Tr}(M+\lambda\left[J^{s}\right]^{2})^{k}$
as
$\operatorname{Tr}(M+\lambda\left[s^{2}\text{Id}+2sA+A^{2}\right])^{k},$
where we have introduced the diagonal matrix $A=\text{diag}(0,A_{2},\dots,A_{n})$. Replace the indeterminate variable $\lambda$ with $\sigma/(2s)$ so that this becomes
$h_{k}^{\sigma}=\operatorname{Tr}\left(\left[M+\sigma\left(A+\frac{A^{2}}{2s}\right)\right]+\frac{s\sigma}{2}\text{Id}\right)^{k}.$
Notice that the expression above may be written as
$\operatorname{Tr}\left[M+\sigma\left(A+\frac{A^{2}}{2s}\right)\right]^{k}+\text{ a linear combination of }h^{\sigma}_{k-1},\dots,h^{\sigma}_{2},h^{\sigma}_{1},h^{\sigma}_{0}.$
Therefore, the Manakov integrals $h_{k}^{\sigma}$ are functionally equivalent to the collection
$\operatorname{Tr}\left[M+\sigma\left(A+\frac{A^{2}}{2s}\right)\right]^{k}$
for $k=1,2,\dots,n$. These integrals have a well-defined limit as $s\rightarrow\infty$ given by
$g^{\sigma}_{k}=\operatorname{Tr}(M+\sigma A)^{k}$
for arbitrary $\sigma$. In the next section we will verify that these are indeed first integrals of the sub-Riemannian geodesic flow.
3.2. A Lax pair and bi-Hamiltonian structure
The Poisson structure in (3) leads to the equation of motion
(13) $\frac{d}{dt}M=[dH_{\text{sR}},M],$
in $\mathfrak{so}_{+}(n)^{*}$ where the derivative $dH_{\text{sR}}$ is understood to be an element of the Lie algebra. It will be helpful to decompose
$M=\begin{pmatrix}0&-\mu^{T}\\ \mu&\hat{M}\end{pmatrix}$
into the $(n-1)$-dimensional column vector $\mu$ and the skew-symmetric matrix $\hat{M}$ belonging to $\mathfrak{so}(n-1)$. In a slight abuse of notation, by letting $\hat{A}^{-1}$ denote the matrix $\text{diag}(A_{2}^{-1},\dots,A_{n}^{-1})$ the Hamiltonian is conveniently written
$H_{\text{sR}}(M)=\frac{1}{2}\operatorname{Tr}\left(\mu^{T}\hat{A}^{-1}\mu\right)$
from which we find
(14) $dH_{\text{sR}}=\begin{pmatrix}0&-\mu^{T}\hat{A}^{-1}\\ \hat{A}^{-1}\mu&0\end{pmatrix}.$
Together with (13) this allows us to write the equations of motion as
$\frac{d}{dt}\mu$ $=-\hat{M}\hat{A}^{-1}\mu,$
$\frac{d}{dt}\hat{M}$ $=[\mu\mu^{T},\hat{A}^{-1}].$
This admits a Lax pair formulation,
(15) $\frac{d}{dt}(M+\sigma A)=[dH_{\text{sR}}+\sigma B,M+\sigma A],$
for $\sigma$ an arbitrary parameter and $B=\text{diag}(0,1,\dots,1)$. It follows that $M+\sigma A$ evolves within the same conjugacy class. The traces of $(M+\sigma A)^{k}$ are therefore integrals of motion. Since $\sigma$ is arbitrary, the coefficients of $\sigma$ in the expansion
(16) $g^{\sigma}_{k}=\operatorname{Tr}(M+\sigma A)^{k}=\sum_{l=0}^{k}\sigma^{k-l}g_{k,l}(M)$
are first integrals, which we name the sub-Riemannian Manakov integrals, and denote by $g_{k,l}$.
Remark 3.1.
It is interesting to note that the sub-Riemannian Manakov integrals are the same as the ordinary Manakov integrals in (5) for a ‘flat’ rigid body whose diagonal mass matrix satisfies $J^{2}=A$. The body is flat because it has no extent in the first principal direction since $J_{1}=0$. The equations of motion for the flat body are, however, not the same as those for the sub-Riemannian problem. This can be seen by noting that the inertia operator $\mathbb{J}$ for a flat rigid body (with only $J_{1}$ equal to zero) is invertible, yet the cometric $M\mapsto dH_{\text{sR}}$ for the sub-Riemannian metric is not invertible.
We now turn our attention to establishing a bi-Hamiltonian structure for the equations of motion. For any symmetric matrix $S$ there is an alternative Lie bracket on $\mathfrak{so}(n)$ given by
$[X,Y]_{S}=XSY-YSX.$
We can use this to modify (3) to produce an alternative Poisson structure on $\mathfrak{so}(n)^{*}$ given by
$\{f,g\}_{S}=\langle M,[df,dg]_{S}\rangle.$
For a given Hamiltonian $H$ suppose that there is another function $F$ for which
$\{\varphi,H\}=\{\varphi,F\}_{S}$
holds for all $\varphi$. The Hamiltonian system generated by $H$ is therefore equivalent to the system generated by $F$ with respect to the altered Poisson structure. If $S$ is not a scalar multiple of the identity then we say that this system is bi-Hamiltonian.
Proposition 3.1.
The sub-Riemannian Hamiltonian system generated by $H_{\text{sR}}$ on $\mathfrak{so}_{+}(n)^{*}$ is bi-Hamiltonian. Specifically, the system is equivalent to the Hamiltonian system generated by
(17) $F(M)=\operatorname{Tr}\left(\mu^{T}\hat{A}^{-1}\mu\right)+\frac{1}{2}\operatorname{Tr}\left(\hat{M}^{T}\hat{A}^{-1}\hat{M}\hat{A}^{-1}\right)$
with respect to the altered Poisson structure {,}$\{\leavevmode\nobreak\ ,\leavevmode\nobreak\ \}_{A}$.
Proof.
From the definitions of the Poisson brackets, we must show that
(18) $\langle M,[X,dH_{\text{sR}}]\rangle=\langle M,[X,dF]_{A}\rangle$
for all $X$ in $\mathfrak{so}(n)$. As we are identifying $\mathfrak{so}(n)$ with its dual via the trace form, one can show that
$\langle M,[X,dH_{\text{sR}}]\rangle=\langle X,[dH_{\text{sR}},M],\rangle\text{and }\langle M,[X,dF]_{A}\rangle=\langle X,[A,M]_{dF}\rangle.$
Therefore, it suffices to show
(19) $\langle X,[dH_{\text{sR}},M]\rangle=\langle X,[A,M]_{dF}\rangle.$
The derivative of $F$ is found to be
(20) $dF=\begin{pmatrix}0&-\mu^{T}\hat{A}^{-2}\\ \hat{A}^{-2}\mu&\hat{A}^{-1}\hat{M}\hat{A}^{-1}\end{pmatrix}.$
which when combined with (14) allows us to compute and verify the important identity $[dH_{\text{sR}},M]=[A,M]_{dF}$, from which (19) follows.
Remark 3.2.
The bi-Hamiltonian structure we have constructed is a variation of that for the rigid body given in [27] (see also the construction in [5]) who are careful to note that if any of the $J_{i}$ are zero then this “…can also be employed for flat bodies, but the construction of the Hamiltonian functions is more complicated”. Indeed, we observe that the altered Hamiltonian $F$ in (17) differs from the form given in [27, Eq. 3.8]. Nevertheless, as confirmed in their concluding remarks, the proof that the integrals are in involution using a recursion relation may still be applied. Indeed, one can reproduce their arguments exactly to show that the sub-Riemannian Manakov integrals commute.
3.3. Proof of Theorem 1
The integrability of the sub-Riemannian geodesic system follows directly from an application of Mishchenko and Fomenko’s results which generalise the Euler equations of the rigid body to arbitrary semisimple Lie algebras.
Let $\mathfrak{g}$ be the complex semisimple Lie algebra $\mathfrak{gl}_{n}\mathbf{C}$ with ‘compact form’ $\mathfrak{g}_{u}=\mathfrak{u}(n)$ and the ‘normal form’ $\mathfrak{g}_{n}=\mathfrak{so}(n)$. One may verify the identity
(21) $[dH_{\text{sR}},A]=[M,B],$
from which the integrability of the equations of motion (13) follow directly from [24, Theorem 5.2]. More specifically, the integrals may be chosen from the collection of sub-Riemannian Manakov integrals, which are all of the form
$f(M+\sigma A),$
for $f$ a Casimir. In order to apply this theorem we note that although $A$ has a single zero-eigenvalue, it is still a regular element in $\mathfrak{gl}_{n}\mathbf{C}$ since all of its eigenvalues are distinct. The Lax pair formulation (15) also follows (see for instance the statement in [9, Theorem 1]).
Remark 3.3.
There is a straightforward generalisation which replaces the distribution of tangent vectors in (1) with block matrices
$\Omega=\begin{pmatrix}0&-\omega^{T}\\ \omega&0\end{pmatrix},$
where $\omega$ is a $k\times(n-k)$-matrix, with sub-Riemannian length
$\sqrt{\operatorname{Tr}(\omega^{T}\hat{A}\omega)},$
for $\hat{A}=\text{diag}(A_{n-k},\dots,A_{n})$. When interpreted as a rolling distribution this corresponds to the problem of rolling the (oriented) Grassmannian $\mathbb{Gr}_{k,n}^{+}$ along its affine tangent space. This problem has been considered extensively before in [17, 19] (see also the references therein) for the particular case of the standard Euclidean metric where $\hat{A}$ may be taken to be the idenity matrix. We remark however, that the result of Mishchenko and Fomenko used above does not work for $k>1$. The reason for this is that, although Eq. (21) holds for $A=\textrm{diag}(0,\dots,0,A_{n-k},\dots,A_{n})$ and $B=\textrm{diag}(0,\dots,0,1,\dots,1)$, the matrix $A$ is not regular, and so the proof in [24] no longer applies.
4. Further Work
A curious feature which distinguishes sub-Riemannian geometry from ordinary Riemannian geometry is the existence of singular geodesics. These are length minimising curves which, unlike normal geodesics, are not the projections of integral curves in the cotangent bundle for the Hamiltonian flow. In other words, these are geodesics that do not arise from solutions of the equations of motion, and are typically more difficult to find. Therefore, it is natural to ask what, if any, singular geodesics are present for the sub-Riemannian metric on $\mathbf{SO}(n)$? We note that for $n=3$ the distribution is contact and so all geodesics are normal [1, Proposition 4.38].
The relative equilibria of the standard Euler equations on $\mathfrak{so}(3)$ correspond to the well-known steady motions of a rigid body around its principal axes. In our example for $n=3$, the analogous solutions can be seen from Figure 1 to correspond to motions where the ball rolls in a straight line across the table. For $n>3$ a classification of all such relative equilibria is more challenging, and corresponds to finding critical points of the sub-Riemannian Hamiltonian on the adjoint orbits. This task has been considered for the case of geodesic flows in [31].
It is somewhat frustrating that we must assume the $A_{2},\dots,A_{n}$ to be distinct. This frustration is shared with the rigid body, where the integrability results of Manakov and Mishchenko & Fomenko both require the $J_{i}$ to be distinct. In the event that these become equal we obtain the so-called symmetric body. The proof of integrability breaks down as dependencies form between the Manakov integrals which were previously independent. On the other hand, the added symmetry of the body gives rise to additional conserved momenta. These, however, do not necessarily commute with one another, so one must consider the related notion of non-commutative integrability, as performed in [14, 15]. One could adapt their work to the sub-Riemannian case where some of the $A_{i}$s become equal.
Finally, we conclude with a conjecture. The famous Mishchenko-Fomenko conjecture (now a theorem [32], see also [8]) asked whether every Lie group admits an integrable Riemannian metric. This was intially proven by Mishchenko & Fomenko for semisimple Lie groups in [24]. Our work motivates a related sub-Riemannian extension of this question: for any (semisimple) Lie group $G$, does their exist an integrable sub-Riemannian metric of any given rank $k$? The known Riemannian case corresponds to $k=\dim G$. In this paper we have showed that for $G=\mathbf{SO}(n)$, the result is true for the distribution $\mathcal{D}$ with rank $k=n-1$. The general case would shed light on the class of systems discussed in [7] and [21].
Acknowledgements
We wish to express our gratitude to the mathematics department at the University of Michigan for enabling this collaboration. We also thank Richard Montgomery for his helpful and constructive correspondence throughout the preparation of this paper. This research was supported in part by NSF grant 2103026 and AFOSR grants FA9550-23-1-0215 and FA9550-23-1-0400